Find an equation of the line having the specified slope and containing the indicated point. Write your answer in slope-intercept form.
m = -3; (- 7, 6)

Answer by checkley77(12844)

Y=mX+b
m = -3; (- 7, 6)
6=-3*-7+b
6=21+b
b=6-21
b=-15 ans. for the Y intercept.
Y=-X/3-15 ans.